OpenStudy (anonymous):

int_{}^{}10^x dx = integrate it. if can show me the step please.

OpenStudy (amistre64):

since your missing the ln(10) id find a suitable "1" to use on it

OpenStudy (anonymous):

the value not needed. i just need the way to solve it. the notes given by my lecturer is incomplete.=(

OpenStudy (amistre64):

10^x comes from a: 10^x but when we derivative it we get: 10^x ln(10) the ln(10) is just a constant so we can borrow it from a "1". ln(10)/ln(10) is a good "1" to use

OpenStudy (amistre64):

its either that or fill you mind with all the u sub stuff for something so generic :)

OpenStudy (amistre64):

u = 10^x du = 10^x ln(10) dx du/ln(10) = 10^x dx int: du/ln(10)

OpenStudy (amistre64):

\[\int 10^x\ dx\] \[\int 1*10^x\ dx\] \[\int \frac{ln(10)}{ln(10)}*10^x\ dx\] \[\frac{1}{ln(10)}\int ln(10)*10^x\ dx\]

OpenStudy (anonymous):

the 10^x dx need to sub it with dy/ln10 ?

OpenStudy (amistre64):

need to? no. Can you? sure ...

OpenStudy (anonymous):

then the final answer is??

OpenStudy (amistre64):

\[\int 10^x\ dx\] \begin{array}l u=10^x\\ du=10^x\ ln(10)\ dx\\ \frac{1}{ln(10)}du=10^x\ dx \end{array} \[\int\frac{1}{ln(10)}du\] \[\frac{1}{ln(10)}\int\ du\]

OpenStudy (amistre64):

the final answer is what you get to when you finish off the integration :)

OpenStudy (anonymous):

is the final answer 1/ln(10) * 10^x ?

OpenStudy (amistre64):

yep, sure is; +C if your doing the indefinite integral

OpenStudy (anonymous):

ok... so if i did not sub the with the 1/ ln(10), would the answer still the same?

OpenStudy (jamesj):

It must be the same.

OpenStudy (amistre64):

there is only one answer ....

OpenStudy (amistre64):

there are many methods to get there; but only one answer

OpenStudy (anonymous):

ok...i understand it now.thanks a lot.=)

OpenStudy (amistre64):

youre welcome

OpenStudy (jamesj):

You can generalize this problem. It is true for all a > 0 that \[ \frac{d \ }{dx} a^x = \ln a \ . a^x \] Hence \[ \int a^x \ dx = \frac{1}{\ln a} a^x \ + C \]

OpenStudy (jamesj):

Notice when a = e, than ln a = ln e = 1, hence \[ \int e^x = e^x + C \] which is what we expect.